OpenStudy (anonymous):

Solve 1n 2 + 1n x = 5. Round to the nearest thousandth, if necessary.

OpenStudy (kropot72):

\[\ln 2+\ln x=\ln 2x\] \[\ln 2x=5\] Therefore: \[e ^{5}=2x\] Now you can solve for x. Can you do that?

OpenStudy (anonymous):

I got 74.206 ? Correct ?

OpenStudy (kropot72):

@kimblack Your answer is correct :)